A parallelepiped is defined by the following vectors |π‘Ž| = (2, 3, βˆ’1), |𝑏| = (βˆ’1, 0, 2),

and |𝑐| = (3, βˆ’1, 2). Find the volume of the parallelepiped.Β 


1
Expert's answer
2021-05-12T05:04:23-0400

The volume is given by the formula "|a\cdot |b\times c||." We can write "b= -\hat{i}+2\hat{k}" and "c=3\hat{i}-\hat{j}+2\hat{k}." "b\times c=\hat{i}(-2\times -1)+\hat{j}(2\times 3-(-1)\times 2)+\hat{k}(-1\times -1)" "=2\hat{i}+8\hat{j}+\hat{k}." Hence "a\cdot(b\times c)=" "(2\hat{i}+3\hat{j}-\hat{k})\cdot (2\hat{i}+8\hat{j}+\hat{k})" = "2\times(2)+(3)\times 8-(1\times 1)=27." Hence volume ="|27|=27" in cube units.